OpenStudy (anonymous):

y varies inversely as twice x. When x=9, y=2. Find y when x = 3. 4 6 9 12 I think its 12.

OpenStudy (mimi_x3):

I think that its 6 idk

OpenStudy (katrinakaif):

The answer is 6..

OpenStudy (anonymous):

For curiosity sake, can you show me... Thanks!

OpenStudy (mimi_x3):

Well, since x= 9 , y=2 then when x = 3 , y = 6 because you times it by 2 lols , does it make sense xD

OpenStudy (anonymous):

but 2*2 = 4 not 9...

OpenStudy (katrinakaif):

Equation for inverse proportions : y = k/x When it says varies inversely as twice x, the equation is y = k/2x Now plug in the values for y and x as y = 2 and y = 9 and you should get that k = 36 Now subsitute in for y = 36/2(3) then y = 6

OpenStudy (anonymous):

I see! Great! Thank you!!!

OpenStudy (katrinakaif):

Srry laaag.. >.<

OpenStudy (anonymous):

Thats okay! As long as I understand, I'm happy :)

OpenStudy (mimi_x3):

LOL, my one was only a guess xD I guessed it correct xD

OpenStudy (katrinakaif):

haha =)